an aquaous solution of 8 molar and 9 molal and molar mass of solute is 50 gm find density
 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
A very simple formula is enough for solving this type of questions. you should memorize this one . Formula is 1/m = d/M - Ms/1000 
Here, m is molality 
M is molarity 
d is density of solution 
Ms is molar mass of solute .
Given, m = 9, M = 8 , Ms = 50 g/mol 
∴ 1/9 = d/8 - 50/1000
⇒ 0.11 = d/8 - 0.05
⇒0.11 + 0.05 = d/8
⇒0.16 = d/8 
⇒d = 0.16 × 8 = 1.28 g/ml 
Hence, density of solution is 1.28 g/ml